# G2P: Gaussian-to-Point Attribute Alignment for Boundary-Aware 3D Semantic Segmentation

### v1

@@ 4.2 Quantitative Comparison
We compare G2P with geometric methods and 2D-3D fusion approaches on the ScanNet v2 validation set. Tab. 1 reports the results. G2P exceeds PT v3 by +1.4 mIoU and BFANet [bfanet] by +1.1 under identical settings. G2P outperforms all geometric and 2D–3D fusion baselines without external pre-training. For reference, UniPre3D [unipre3d] and ODIN [odin] both use external pre-training beyond ScanNet v2.
@@
MinkUNet [Minkwski] CVPR’19 25.3 46.3 15.4 10.6
OctFormer [oct] TOG’23 32.6 53.9 26.5 13.1
CeCo [ceco] CVPR’23 34.0 55.1 24.7 18.1
PonderV2 [ponderv2] CVPR’24 34.6 55.2 27.0 17.5
BFANet [bfanet] CVPR’25 36.0 55.3 29.3 19.3
@@
Results on ScanNet200. We further evaluate on the ScanNet200 [scannet200] validation set, which expands the label space from 20 to 200 fine-grained categories (Tab. 3). G2P attains 36.6 mIoU, 46.0 mAcc, and 83.8 OA, improving over our reproduced PT v3 [PointTFv3] baseline by +2.1 mIoU and remaining competitive with UniPre3D [unipre3d]. We do not report BFANet [bfanet] on ScanNet200 due to the lack of official implementation for this dataset.
@@ 4.3 Qualitative Comparisons
Qualitative results on challenging classes. Fig. 4 compares G2P with PT v3 [PointTFv3] and BFANet [bfanet] on challenging ScanNet v2 classes. Baselines exhibit two primary failure modes, highlighted by yellow boxes. Boxes with stars ($\star$) mark category confusion: in columns (a) door and (e) refrigerator, the baselines misclassify the objects as wall or cabinet due to coplanar geometry and similar textures. G2P correctly segments these objects. Boxes without stars mark incomplete segmentation: in columns (b) window, (c) picture, (d) curtain, and (f) shower curtain, baselines fail to cover the full extent of thin structures. G2P produces sharper, more complete masks for these objects. These trends align with Tab. 2, where G2P achieves substantial gains on refrigerator (improvements of 6.0% in IoU vs. PT v3), confirming that G2P effectively mitigates both failure modes.
@@ 4.4 Ablation Studies
Results on Additional Benchmarks. To further validate G2P’s effectiveness, we evaluate on the ScanNet++ [scannetpp] and Matterport3D [matterport3d] validation sets. As shown in Tab. 6, G2P generalizes well across datasets, achieving 48.7 and 55.9 mIoU on ScanNet++ and Matterport3D, outperforming PT v3 by +0.8 and +0.4.
@@
Table 7: Ablation on strategies for Gaussian primitive (opacity) Learning. We compare: (i) Raw 3D Gaussians, which use Gaussian primitives as direct training inputs; (ii) Points with Opacity, which also provide point features and their Gaussian opacity directly to the model; and (iii) G2P (Ours), which distills Gaussian appearance to the segmentation model. All models are evaluated on ScanNet v2 validation set.
@@
MinkUNet [Minkwski] 39.2M 71ms 1.6G 29ms 1.4G 72.3
OctFormer [oct] 44.0M 259ms 4.2G 94ms 4.4G 74.3
PT v3 [PointTFv3] 46.2M 132ms 5.6G 79ms 1.9G 77.0
@@ 5 Conclusion
We presented G2P, a novel approach for point cloud semantic segmentation that integrates 3D Gaussian Splatting to unify geometric and appearance cues in 3D space. Our three-component approach, comprising Gaussian-to-point feature augmentation, opacity-guided appearance learning, and scale-based boundary extraction, addresses the fundamental geometric bias. By augmenting point features with Gaussian attributes, G2P enhances semantic segmentation on point clouds, particularly for objects that are geometrically ambiguous but have distinctive appearances. Extensive experiments on standard benchmarks demonstrate state-of-the-art performance and strong generalization across datasets. Furthermore, we conduct ablations to validate the effectiveness of GS-guided feature augmentation, establishing its practical value for 3D scene understanding.

### v2

@@ 4.2 Quantitative Comparison
We compare G2P with geometric methods and 2D-3D fusion approaches on the ScanNet v2 validation set. Tab. 1 reports the results. G2P exceeds PT v3 by +1.4 mIoU and BFANet [63] by +1.1 under identical settings. G2P outperforms all geometric and 2D–3D fusion baselines without external pre-training. For reference, UniPre3D [49] and ODIN [20] both use external pre-training beyond ScanNet v2.
@@
MinkUNet [7] CVPR’19 25.3 46.3 15.4 10.6
OctFormer [48] TOG’23 32.6 53.9 26.5 13.1
CeCo [64] CVPR’23 34.0 55.1 24.7 18.1
PonderV2 [66] CVPR’24 34.6 55.2 27.0 17.5
BFANet [63] CVPR’25 36.0 55.3 29.3 19.3
@@
Results on ScanNet200. We further evaluate on the ScanNet200 [42] validation set, which expands the label space from 20 to 200 fine-grained categories (Tab. 3). G2P attains 36.6 mIoU, 46.0 mAcc, and 83.8 OA, improving over our reproduced PT v3 [53] baseline by +2.1 mIoU and remaining competitive with UniPre3D [49]. We do not report BFANet [63] on ScanNet200 due to the lack of official implementation for this dataset.
@@ 4.3 Qualitative Comparisons
Qualitative results on challenging classes. Fig. 4 compares G2P with PT v3 [53] and BFANet [63] on challenging ScanNet v2 classes. Baselines exhibit two primary failure modes, highlighted by yellow boxes. Boxes with stars ($\star$) mark category confusion: in columns (a) door and (e) refrigerator, the baselines misclassify the objects as wall or cabinet due to coplanar geometry and similar textures. G2P correctly segments these objects. Boxes without stars mark incomplete segmentation: in columns (b) window, (c) picture, (d) curtain, and (f) shower curtain, baselines fail to cover the full extent of thin structures. G2P produces sharper, more complete masks for these objects. These trends align with Tab. 2, where G2P achieves substantial gains on refrigerator (improvements of 6.0% in IoU vs. PT v3), confirming that G2P effectively mitigates both failure modes.
@@ 4.4 Ablation Studies
Results on Additional Benchmarks. To further validate G2P’s effectiveness, we evaluate on the ScanNet++ [59] and Matterport3D [4] validation sets. As shown in Tab. 6, G2P generalizes well across datasets, achieving 48.7 and 55.9 mIoU on ScanNet++ and Matterport3D, outperforming PT v3 by +0.8 and +0.4.
@@
Table 7: Ablation on strategies for Gaussian primitive (opacity) Learning. We compare: (i) Raw 3D Gaussians, which use Gaussian primitives as direct training inputs; (ii) Points with Opacity, which also provide point features and their Gaussian opacity directly to the model; and (iii) G2P (Ours), which distills Gaussian appearance to the segmentation model. Here, $c^{\prime}$ and $n^{\prime}$ denote Gaussian color and normal features aggregated from neighboring points based on Euclidean distance.
@@
MinkUNet [7] 39.2M 71ms 1.6G 29ms 1.4G 72.3
OctFormer [48] 44.0M 259ms 4.2G 94ms 4.4G 74.3
PT v3 [53] 46.2M 132ms 5.6G 79ms 1.9G 77.0
@@ 5 Conclusion
We presented G2P, a novel approach for point cloud semantic segmentation that integrates 3DGS to unify geometric and appearance cues in 3D space. Our three-component approach, comprising Gaussian-to-point feature augmentation, opacity-guided appearance learning, and scale-based boundary extraction, addresses the fundamental geometric bias. By augmenting point features with Gaussian attributes, G2P enhances semantic segmentation on point clouds, particularly for objects that are geometrically ambiguous but have distinctive appearances. Extensive experiments on standard benchmarks demonstrate state-of-the-art performance and strong generalization across datasets. Furthermore, we conduct ablations to validate the effectiveness of GS-guided feature augmentation, establishing its practical value for 3D scene understanding.